# Corks, exotic 4-manifolds and genus functions

### v3

@@ 3.3. Modifying 2-handlebodies into exotic Stein manifolds
Let $X$ be a 4-dimensional 2-handlebody, and let $K$ be the attaching circle of a 2-handle of $X$. We consider a handlebody diagram of $X$. Akbulut and the author [11] introduced two local diagrammatic operations shown in Figure 1. For a positive integer $p$, the left side operation is called a $W^{-}(p)$-modification to $K$, and the right side operation is called a $W^{+}(p)$-modification to $K$. We note that these operations do not change the framing coefficient $m$ of $K$. When we do not specify the coefficient $p$, these operations are called a $W^{-}$-modification and a $W^{+}$-modification, respectively.
@@
Clearly, a $W^{-}$-modification attaches an algebraically canceling pair of 1- and 2-handles to $X$. Thus, this operation is an attachment of an invertible homology cobordism to $X$, and its inverse cobordism is its upside down cobordism. Furthermore, the homotopy extension property (see [39, Proposition 0.1.8]) implies that this invertible cobordism has a deformation retraction onto $\partial X$. Consequently, the result of any $W^{-}$-modification and hence also any $W^{+}$-modification topologically resembles $X$. Furthermore, the result of any $W^{-}$-modification admits a deformation retraction onto $X$. We note that the 4-manifold obtained by a $W^{+}$-modification becomes diffeomorphic to the original 4-manifold by attaching the upside down cobordism of the invertible cobordism given by the corresponding $W^{-}$-modification ([11, Proposition 4.5]).
By drawing Legendrian pictures, we obtain Legendrian versions of $W^{-}$- and $W^{+}$-modifications for Legendrian handlebodies, and these have the following two effects ([11, Proposition 4.7]). (1) Each of these two operations modifies a Stein handlebody into another Stein handlebody. (2) If we apply a $W^{+}$-modification to a Legendrian knot $K$ of a Legendrian handlebody, then the Thurston–Bennequin number of the result of $K$ increases by $p$, and the Thurston–Bennequin number for the attaching circle of the 2-handle of $W$ is $+2$.
Thus, by applying $W^{+}$-modifications to a 2-handlebody $X$ and adding zig-zags to each Legendrian knot (i.e., stabilizations), we obtain a Stein handlebody, implying the following.
@@
For any $4$-dimensional $2$-handlebody $X$, there exists a $4$-dimensional $2$-handlebody $X^{\prime}$ having the following properties.
@@
  • $X^{\prime}$ topologically resembles $X$ and is homeomorphic to a $4$-manifold that admits a deformation retraction onto $X$.
@@
Applying $W^{-}$-modifications several times and replacing some of them with $W^{+}$-operations by cork twists, we obtain arbitrarily many pairwise homeomorphic 2-hanldebodies. In the case where $H_{2}(X)$ is a non-trivial group, by carefully choosing coefficients of $W^{+}$- and $W^{-}$-modifications and applying adjunction inequalities, Akbulut and the author showed that the resulting 2-handlebodies have pairwise algebraically inequivalent genus functions for any orientations ([11, Lemma 5.13 and Proposition 6.2]), giving the following result.
@@
For any $4$-dimensional $2$-handlebody $X$ with a non-trivial second homology group and any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{0},X_{1},\dots,X_{n}$ having the following properties.
  • Each $X_{i}$ topologically resembles $X$ and admits an embedding into $X$.
  • $X_{0}$ admits a deformation retraction onto $X$ and is obtained from $X$ by attaching an invertible homology cobordism.
@@ 3.4. Modifying generalized corks into exotic 4-manifolds
In this subsection, we review the method of Akbulut and Ruberman [8] that modifies a given generalized cork into exotic 4-manifolds. We also give minor improvements of their results.
@@
There exist a closed $3$-manifod $N$ and an invertible homology cobordism $\mathbb{P}$ from $M$ to $N$ having the following properties.
@@
Let $f_{1},f_{2}$ be self-diffeomorphisms of $M$ that preserve the orientations, and let $X_{1},X_{2}$ be $4$-manifolds with $\partial X_{1}=\partial X_{2}=M$. If $f_{2}^{-1}\circ f_{1}$ does not extend to any diffeomorphism $X_{1}\to X_{2}$, then the following hold.
@@
  • Each $X_{i}\cup_{f_{i}}\mathbb{P}$ admits a deformation retraction onto $X_{i}$ and topologically resembles $X_{i}$.
@@
Akbulut and Ruberman constructed a cobordism $\mathbb{P}$ in Theorem 3.5 using an arbitrary doubly slice knot in $S^{3}$ having certain properties, and they showed that the specific knot $J$ in $S^{3}$ given by [8, Figure 2] can be used for the construction. In the above theorem, we use the cobordism $\mathbb{P}$ corresponding to the knot $J$ in order to guarantee the homotopy equivalence. The rest of the properties hold for any $\mathbb{P}$ in Theorem 3.5 as seen from our proof.
Proof.
@@
We set $X^{\prime}_{i}=X\cup_{f_{i}}\mathbb{P}$ for each $i$. Then each $X^{\prime}_{i}$ deformation retracts to $X$ by Lemma 3.8 which will be given later. Since $\mathbb{P}$ is a homology cobordism, we see that $X^{\prime}_{i}$ topologically resembles $X_{i}$.
@@
Composing the last diffeomorphisms and $\Phi$, we obtain a diffeomorphism $X_{1}\cup_{\mathrm{id}_{M}}(I\times M)\to X_{2}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{2}^{-1}\circ f_{1}$. Since each $X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ is diffeomorphic to $X_{i}$ fixing the boundary $M$ pointwise, it follows that $f_{2}^{-1}\circ f_{1}$ extends to a diffeomorphism $X_{1}\to X_{2}$, giving a contradiction. Therefore, $X^{\prime}_{1}$ is not diffeomorphic to $X^{\prime}_{2}$. ∎
The homotopy equivalence of $X_{i}\cup_{f_{i}}\mathbb{P}$ and $X_{i}$ is claimed in [8, Theorem A], but the proof is stated only in the case where $X_{i}$ is contractible. For completeness, we give a proof of the general case for $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) by showing the lemma below.
Lemma 3.8 (cf. [8]).
@@
In order to construct $\mathbb{P}$, Akbulut and Ruberman [8] used a certain link $L$ in $M$ and a certain concordance $C(\subset I\times S^{3})$ from the unknot $U$ to the knot $J$ in [8, Figure 2]. Let $\nu(L)$ and $\nu(C)$ respectively denote the tubular neighborhoods of $L$ and $C$, and let $n$ denote the number of the components of $L$.
They constructed $\mathbb{P}$ from $I\times(M-\textnormal{int}\,\nu(L))$ by attaching $n$ copies of $(I\times S^{3})-\textnormal{int}\,\nu(C)$ along $I\times\partial\nu(L)$ so that the resulting negative boundary remains $M$. Note that there is a diffeomorphism from $[-1,0]\times M\cup_{0\times(M-\textnormal{int}\,\nu(L))}I\times(M-\textnormal{int}\,\nu(L))$ to $I\times M$ that maps $(0\times\nu(L))\cup(I\times\partial\nu(L))$ to $1\times\nu(L)$. Thus, $\mathbb{P}$ is obtained from $I\times M$ by attaching $n$ copies of the closure of $(I\times S^{3})-\nu(C)$ using a diffeomorphism from each component of $1\times\nu(L)$ to $\nu(m)$, where $m$ denotes the meridian of the unknot $U$ in $0\times S^{3}$, and $\nu(m)$ denotes a tubular neighborhood of $m$ in $\partial(I\times S^{3}-\textnormal{int}\,\nu(C))$.
Akbulut and Ruberman gave a handlebody diagram of the exterior of a slice disk for $J$ in $D^{4}$ ([8, Figure 4.(a)]), and it is easy to see that this 4-manifold is diffeomorphic to the closure of $(I\times S^{3})-\nu(C)$. Since the attaching circle of the unique 2-handle of this handlebody is homotopic to a meridian of the dotted circle, the closure of $(I\times S^{3})-\nu(C)$ deformation retracts to the tubular neighborhood $\nu(m)$. Therefore, $\mathbb{P}$ admits a deformation retraction onto $0\times M$. ∎
@@
(1) The inverse cobordism of $\mathbb{P}$ used in Theorem 3.6 (see Remark 3.7) is its upside down cobordism $\overline{\mathbb{P}}$, though we do not use this fact in this section. This claim is not stated in [8], but it is implicit and follows from their claim that the knot $J$ in [8, Figure 2] is superslice, that is, the double of a slice disk for $J$ is an unknotted 2-sphere in $S^{4}$ (see the proof of [8, Proposition 2.6]). The supersliceness can be alternatively proved using [51, Proposition 4.3], and it is clear that the slice disk obtained in [8] gives a concodance in $I\times S^{3}$ from the unknot to $J$ such that the double is isotopic to the product concordance fixing $\{0,1\}\times S^{3}$ pointwise. Therefore, it follows from their construction of $\mathbb{P}$ that the double $\mathbb{P}\cup_{\mathrm{id}_{M}}\overline{\mathbb{P}}$ admits a diffeomorphism to the product $I\times M$ whose restriction to the boundary $\overline{M}\sqcup M$ is the identity map.
(2) The cobordism $\mathbb{P}$ is obtained from $I\times M$ by attaching algebraically canceling pairs of 1- and 2-handles, and Lemma 3.8 follows from this fact as well. This fact is implicit in [8, Example 4.1] and can be proved using Akbulut’s cylinder method [4, Section 3.2] as follows. The proof of Lemma 3.8 implies that the cobordism $\mathbb{P}$ is obtained from the disjoint union of $I\times M$ and copies of the exterior of the slice disk for $J$ by attaching copies of the cylinder $I\times S^{1}\times D^{2}$ along $\partial I\times S^{1}\times D^{2}$. Since this cylinder is a union of a 1-handle and a 2-handle, the cobordism $\mathbb{P}$ is obtained from the boundary sum of $I\times M$ with the copies of the exterior by attaching 2-handles. Due to the handlebody diagram of the exterior ([8, Figure 4.(a)]), we can prove the above fact by showing that the attaching circle of the 2-handle in each cylinder cancels a 1-handle of the exterior. We remark that this argument gives a method for drawing handlebody diagrams of $\mathbb{P}$ and each glued 4-manifold $X_{i}\cup_{f_{i}}\mathbb{P}$, once the link $L$ in $M$ is specified.
The corollary below is straightforward from Theorem 3.6.
@@
Let $\{(C,f)\}_{f\in A}$ be a family of pairwise inequivalent generalized corks with $\partial C=M$. Then there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{M}\}}$ of pairwise exotic $4$-manifolds topologically resembling $C$ such that each $C_{f}$ admits a deformation retraction onto $C$.
We note that for any generalized cork $(C,f)$ of infinite order, $\{(C,f^{k})\}_{k\in\mathbb{Z}-\{0\}}$ satisfies the assumption of this corollary. Also, the assumption holds for families given by finite order generalized corks.
@@ 3.5. Infinite order corks
In this subsection, we briefly review Gompf’s infinite order corks ([36]). For integers $r,s,m>0$, let $C(r,s;m)$ be the 4-manifold in Figure 2. (In this notation, we changed the sign of $m$ from the original one in [36] so that we may assume $m>0$.) It is easy to see that $C(r,s;m)$ is a contractible 4-manifold of Mazur type, that is, it admits a handle decomposition consisting of a 0-handle and an algebraically canceling pair of 1- and 2-handles.
@@
Gompf showed that the pair $(C(r,s;m),\tau)$ is an infinite order cork, where the self-diffeomorphism $\tau$ on the boundary is the torus twist on a certain embedded torus. Roughly speaking, a torus twist is (Dehn twist on an $(\textnormal{Dehn twist on an annulus})\times\mathrm{id}_{S^{1}}$, and hence $\tau$ is the identity map outside the tubular neighborhood of the torus. For the precise definition of a torus twist, see [37, Definition 2.1]. We recall an important property of Gompf’s corks. Let $E$ be the 4-manifold given in Figure 3. We denote the obvious torus $T^{2}\times 0\subset T^{2}\times D^{2}$ in the figure by $T$. Also, for an integer $k$, we denote the twist knot in Figure 4 by $T_{k}$. Let $E_{k}$ be the 4-manifold obtained from $E$ by applying the Fintushel–Stern knot surgery ([31]) along the torus $T$ using the knot $T_{k}$.
@@
Gompf [36] proved the following theorem to show that each $(C(r,s;m),\tau)$ is an infinite order cork. For details, see [37, the proof of Theorem 3.5].
@@
Let $r,s,m$ be positive integers, and set $N=r+s+m-3$. Then there exists an embedding of $C(r,s;m)$ into $E\#N\overline{\mathbb{C}P^{2}}$ such that, for each integer $k$, the knot surgered $4$-manifold $E_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the $4$-manifold obtained from $E\#N\overline{\mathbb{C}P^{2}}$ by removing the embedded $C(r,s;m)$ and gluing it back via $\tau^{k}$. Furthermore, this diffeomorphism is isotopic to the identity map on the boundary $\partial E$ of these two $4$-manifolds.

### v5

@@ 3.3. Modifying 2-handlebodies into exotic Stein manifolds
Let $X$ be a 4-dimensional 2-handlebody, and let $K$ be the attaching circle of a 2-handle of $X$. We consider a handlebody diagram of $X$. Akbulut and the author [11] introduced two local diagrammatic operations for a subarc of $K$ shown in Figure 1. For a positive integer $p$, the left side operation is called a $W^{-}(p)$-modification to $K$, and the right side operation is called a $W^{+}(p)$-modification to $K$. We note that these operations do not change the framing coefficient $m$ of $K$. When we do not specify the coefficient $p$, these operations are called a $W^{-}$-modification and a $W^{+}$-modification, respectively.
@@
Every $W^{-}$-modification attaches a homotopically canceling pair of 1- and 2-handles to $X$. Thus, this operation is an attachment of an invertible homology cobordism to $X$, and its upside down cobordism is its inverse cobordism (see Lemma 3.4). Furthermore, this invertible cobordism has a deformation retraction onto $\partial X$ (see Remark 3.5), and hence the result of any $W^{-}$-modification admits a deformation retraction onto $X$. Consequently, the result of any $W^{-}$-modification and hence that of any $W^{+}$-modification is HIHC-equivalent to $X$. In addition, the 4-manifold obtained by any $W^{+}$-modification becomes diffeomorphic to the original 4-manifold $X$ by attaching the upside down cobordism of the invertible cobordism given by the corresponding $W^{-}$-modification ([11, Proposition 4.5]).
By drawing Legendrian pictures, we obtain Legendrian versions of $W^{-}$- and $W^{+}$-modifications for Legendrian handlebodies, and these have the following two effects ([11, Proposition 4.7]). (1) Each of these two operations modifies a Stein handlebody into another Stein handlebody. (2) If we apply a $W^{+}$-modification to a Legendrian knot $K$ of a Legendrian handlebody, then the Thurston–Bennequin number of the result of $K$ increases by $p$, and the Thurston–Bennequin number for the attaching circle of the 2-handle of $W$ is +2+2.
Thus, by applying $W^{+}$-modifications to a Legendrian handlebody and adding zig-zags to each Legendrian knot (i.e., stabilizations), we obtain a Stein handlebody, implying the following.
@@
For any $4$-dimensional 22-handlebody $X$, there exists a $4$-dimensional 22-handlebody $X^{\prime}$ having the following properties.
@@
  • $X^{\prime}$ is HIHC-equivalent to $X$ and is homeomorphic to a $4$-manifold that deformation retracts onto an embedded copy of $X$.
@@
Applying $W^{-}$-modifications several times and replacing some of them with $W^{+}$-operations by cork twists, we obtain arbitrarily many pairwise homeomorphic 2-handlebodies. In the case where $H_{2}(X)\neq 0$, by carefully choosing coefficients of $W^{+}$- and $W^{-}$-modifications and applying adjunction inequalities, Akbulut and the author showed that the resulting 2-handlebodies have pairwise algebraically inequivalent genus functions for any orientations ([11, Lemma 5.13 and Proposition 6.2]), giving the following result.
@@
For any $4$-dimensional 22-handlebody $X$ with $H_{2}(X)\neq 0$ and any positive integer $n$, there exist pairwise exotic $4$-manifolds $X_{0},X_{1},\dots,X_{n}$ having the following properties.
  • Each $X_{i}$ is HIHC-equivalent to $X$ and admits an embedding into $X$.
  • $X_{0}$ is obtained from $X$ by attaching an invertible homology cobordism, and admits a deformation retraction onto $X$.
@@
The proof of this theorem yields a concrete handlebody diagram of each $X_{i}$ from that of $X$ and hence produces various explicit examples of exotic 4-manifolds.
@@ 3.4. Modifying generalized corks into exotic 4-manifolds
In this subsection, we review the method of Akbulut and Ruberman [8] that modifies generalized corks into exotic 4-manifolds. We also improve their results.
@@
There exist a closed $3$-manifod NN and an invertible homology cobordism $\mathbb{P}$ from $M$ to $N$ that satisfy the following.
@@
Let $f_{1},f_{2}$ be self-diffeomorphisms of $M$, and let $X_{1},X_{2}$ be $4$-manifolds with $\partial X_{1}=\partial X_{2}=M$. If $f_{2}^{-1}\circ f_{1}$ does not extend to any diffeomorphism $X_{1}\to X_{2}$, then the following hold.
@@
  • Each $X_{i}\cup_{f_{i}}\mathbb{P}$ is HIHC-equivalent to $X_{i}$ and admits a deformation retraction onto $X_{i}$.
@@
Akbulut and Ruberman constructed a cobordism $\mathbb{P}$ in Theorem 3.8 using an arbitrary doubly slice knot in $S^{3}$ having certain properties, and they showed that the Kinoshita–Terasaka knot $J$ (see [8, Figure 2]) can be used for the construction. In Theorem 3.9, the cobordism $\mathbb{P}$ denotes the one corresponding to the knot $J$ in order to guarantee the homotopy equivalence. The rest of the properties hold for any $\mathbb{P}$ in Theorem 3.8 as seen from our proof.
Proof of Theorem 3.9.
@@
We set $X^{\prime}_{i}=X\cup_{f_{i}}\mathbb{P}$ for each $i$. Then each $X^{\prime}_{i}$ deformation retracts to $X$ by the forthcoming Lemma 3.11. Since $\mathbb{P}$ is a homology cobordism, it follows that $X^{\prime}_{i}$ is HIHC-equivalent to $X_{i}$.
@@
Composing the last diffeomorphisms for $i=1,2$ and $\Phi$, we obtain a diffeomorphism $X_{1}\cup_{\mathrm{id}_{M}}(I\times M)\to X_{2}\cup_{\mathrm{id}_{M}}(I\times M)$ whose restriction to the boundary is $f_{2}^{-1}\circ f_{1}$. Since each $X_{i}\cup_{\mathrm{id}_{M}}(I\times M)$ is diffeomorphic to $X_{i}$ fixing the boundary $M$ pointwise, it follows that $f_{2}^{-1}\circ f_{1}$ extends to a diffeomorphism $X_{1}\to X_{2}$, giving a contradiction. Therefore, $X^{\prime}_{1}$ is not diffeomorphic to $X^{\prime}_{2}$. ∎
The homotopy equivalence of $X_{i}\cup_{f_{i}}\mathbb{P}$ and $X_{i}$ is claimed in [8, Theorem A], but the proof is given only in the case where $X_{i}$ is contractible, and the argument does not seem to extend to the general case. For completeness, we prove the general case for $\mathbb{P}$ used in Theorem 3.9 (see Remark 3.10) by showing the lemma below.
Lemma 3.11.
@@
In order to construct $\mathbb{P}$, Akbulut and Ruberman [8] used a certain link $L$ in $M$ and a certain concordance $C(\subset I\times S^{3})$ from the unknot $U$ to the Kinoshita–Terasaka knot $J$. Let $\nu(L)$ and $\nu(C)$ respectively denote closed tubular neighborhoods of $L$ and $C$, and let $\mathring{\nu}(L)$ and $\mathring{\nu}(C)$ denote the open tubular neighborhoods. Let $n$ be the number of the components of $L$.
They constructed $\mathbb{P}$ from $I\times(M-\mathring{\nu}(L))$ by attaching $n$ copies of $(I\times S^{3})-\mathring{\nu}(C)$ along $I\times\partial\nu(L)$ so that the resulting negative boundary remains $M$. Note that there is a diffeomorphism from $1,0]\times M\cup_{0\times(M-\mathring{\nu}(L))}I\times(M-\mathring{\nu}(L))$ to $I\times M$ that maps $(0\times\nu(L))\cup(I\times\partial\nu(L))$ to $1\times\nu(L)$. Thus, $\mathbb{P}$ is obtained from $I\times M$ by attaching $n$ copies of $(I\times S^{3})-\mathring{\nu}(C)$ using a diffeomorphism from each component of $1\times\nu(L)$ to $\nu(m)$, where $m$ denotes the meridian of the unknot $U$ in $0\times S^{3}$, and $\nu(m)$ denotes a closed tubular neighborhood of $m$ in $\partial(I\times S^{3}-\mathring{\nu}(C))$.
Akbulut and Ruberman gave a handlebody diagram of the exterior of a slice disk for $J$ in $D^{4}$ ([8, Figure 4.(a)]), and it is easy to see that this 4-manifold is diffeomorphic to $(I\times S^{3})-\mathring{\nu}(C)$. Note that $m$ is isotopic to the meridian of a dotted circle in the diagram. Since the unique 2-handle of this handlebody and the 1-handle corresponding to the other dotted circle form a homotopically canceling pair, Remark 3.5 implies that $(I\times S^{3})-\mathring{\nu}(C)$ deformation retracts to the tubular neighborhood $\nu(m)$. Therefore, $\mathbb{P}$ admits a deformation retraction onto the boundary component $0\times M$. ∎
@@
(1) An inverse cobordism of the cobordism $\mathbb{P}$ used in Theorem 3.9 (see Remark 3.10) is its upside down cobordism $\overline{\mathbb{P}}$, though we do not use this fact in this section. This claim is not stated in [8], but it is implicit and follows from their claim that the Kinoshita–Terasaka knot $J$ is superslice, that is, the double of a slice disk for $J$ is an unknotted 2-sphere in $S^{4}$ (see the proof of [8, Proposition 2.6]). The supersliceness can be alternatively proved using [54, Proposition 4.3], and it is clear from the latter proof that the slice disk obtained in [8] gives a concodance in $I\times S^{3}$ from the unknot to $J$ such that the double is isotopic to the product concordance fixing $\{0,1\}\times S^{3}$ pointwise. Therefore, it follows from their construction of $\mathbb{P}$ that the double $\mathbb{P}\cup_{\mathrm{id}_{M}}\overline{\mathbb{P}}$ admits a diffeomorphism to the product $I\times M$ whose restriction to the boundary $\overline{M}\sqcup M$ is the identity map.
(2) The cobordism $\mathbb{P}$ is obtained from $I\times M$ by attaching homotopically canceling pairs of 1- and 2-handles, and Lemma 3.11 and the claims in the above (1) follow from this fact and Lemma 3.4 as well. This fact is implicit from the handlebody diagram of an example of a contractible 4-manifold in [8, Example 4.1] and can be proved using Akbulut’s cylinder method [4, Section 3.2] as follows. The proof of Lemma 3.11 implies that the cobordism $\mathbb{P}$ is obtained from the disjoint union of $I\times M$ and copies of the exterior of the slice disk for $J$ by attaching copies of the cylinder $I\times S^{1}\times D^{2}$ along $\partial I\times S^{1}\times D^{2}$. Since this cylinder is a union of a 1-handle and a 2-handle, the cobordism $\mathbb{P}$ is obtained from the boundary sum of $I\times M$ with the copies of the exterior by attaching 2-handles. Due to the handlebody diagram of the exterior ([8, Figure 4.(a)]), we can prove the aforementioned fact by canceling a 1-handle of each copy of the exterior using the 2-handle of the cylinder. We remark that this argument gives a method for drawing handlebody diagrams of $\mathbb{P}$ and each glued 4-manifold $X_{i}\cup_{f_{i}}\mathbb{P}$, once the link $L$ in $M$ is specified.
We obtain the following corollary, improving a result of Akbulut and Ruberman.
@@
For any family $\{(C,f)\}_{f\in A}$ of pairwise inequivalent generalized corks, there exists a family $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ of pairwise exotic $4$-manifolds which are HIHC-equivalent to $C$.
Proof.
We set $M=\partial C$, and let $\mathbb{P}$ be the cobordism used in Theorem 3.9. For each $f\in A\sqcup\{\mathrm{id}_{\partial C}\}$, let $C_{f}$ denote the 4-manifold $C\cup_{f}\mathbb{P}$. Then, it follows from Theorem 3.9 that $\{C_{f}\}_{f\in A\sqcup\{\mathrm{id}_{\partial C}\}}$ is a desired family. ∎
We note that for any generalized cork $(C,f)$ of infinite order, $\{(C,f^{k})\}_{k\in\mathbb{Z}-\{0\}}$ satisfies the assumption of this corollary. The assumption holds for families given by finite order generalized corks as well.
@@ 3.5. Infinite order corks
In this subsection, we briefly review Gompf’s infinite order corks ([37]). For integers $r,s,m>0$, let $C(r,s;m)$ be the 4-manifold in Figure 2. (In this notation, we changed the sign of $m$ from the original one in [37] so that we may assume $m>0$.) It is easy to see that $C(r,s;m)$ is a contractible 4-manifold of Mazur type, that is, it admits a handle decomposition consisting of a 0-handle and an algebraically canceling pair of 1- and 2-handles (i.e., a 1-handle and a 2-handle whose attaching circle intersects the belt sphere of the 1-handle algebraically once). Note that such a pair attached to the 0-handle is necessarily a homotopically canceling pair. The boundary $\partial C(r,s;m)$ gives infinitely many pairwise non-homeomorphic homology 3-spheres by varying $r,s,m$ ([37, p.2477]).
@@
Gompf proved that each pair $(C(r,s;m),\tau)$ is an infinite order cork, where the self-diffeomorphism $\tau$ on the boundary is the torus twist on a certain embedded torus. Roughly speaking, a torus twist is $(\textnormal{Dehn twist on an annulus})\times\mathrm{id}_{S^{1}}$, and hence $\tau$ is the identity map outside the tubular neighborhood of the torus. For the precise definition of a torus twist, see [38, Definition 2.1].
We recall an important property of Gompf’s corks. Let $E$ be the 4-manifold given in Figure 4. We denote the obvious torus $T^{2}\times 0\subset T^{2}\times D^{2}$ in the figure by $T$. Also, for an integer $k$, we denote the twist knot in Figure 4 by $T_{k}$. Let $E_{k}$ be the 4-manifold obtained from $E$ by applying the Fintushel–Stern knot surgery ([31]) along the torus $T$ using the knot $T_{k}$. Gompf [37] proved the following theorem to show that each $(C(r,s;m),\tau)$ is an infinite order cork. For details, see [38, the proof of Theorem 3.5].
@@
Figure 4. Twist knot $T_{k}$
@@
Let $r,s,m$ be positive integers, and set N=r+s+m3N=r+s+m-3. Then there exists an embedding of $C(r,s;m)$ into $E\#N\overline{\mathbb{C}P^{2}}$ such that for each integer $k$, the knot surgered $4$-manifold $E_{k}\#N\overline{\mathbb{C}P^{2}}$ is diffeomorphic to the $4$-manifold obtained from $E\#N\overline{\mathbb{C}P^{2}}$ by removing the embedded $C(r,s;m)$ and gluing it back via $\tau^{k}$. Furthermore, this diffeomorphism is isotopic to the identity map on the boundary $\partial E$ of these two $4$-manifolds.